[Brahms-l] Shift report 20040215 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 15 2004 - 23:46:03 EST
------------------------------------------------------------
Date: Sun Feb 15 23:46:03 2004, Shift: 16:00-24:00
Supervisor: EJK, with: IA
Log book#15 pages 58 - 58

Shift Summary:
------------------------------------------------------------
1. Beam for 5 hours. At the end of the shift, MCR had a blue QLI.

2. The settings for the entire shift:
  FS at 4 degrees, full field, A polarity
  MRS at 45 degrees, 1050A, A polarity
   Data runs: 10377(started by Djamel),10378, 10379, 10382, 10383
   Pedestal runs: 10380(tpc), 10381(non-tpc)
 Summary:
   ~550k trig2 & ~770k trig3 taken.

3. No problems detected.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 15 23:46:20 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 15 2004 - 23:46:39 EST